# textwidth=0.8textheight=0.9

### v1

@@ 3 Kolmogorov-Arnold Networks
[…]
@@ 3.1 Two-Layer KANs
[…]
@@ 3.2 Multi-Layer KANs
A multi-layer KAN [50] is defined as,
@@
where each $\phi_{k}^{(m)}$ remains a univariate function, but organized in a layered, nested manner. This architecture extends the classical depth-2 superposition model by allowing each univariate component to be represented as a composition of multiple univariate transformations. Consequently, each “outer” function $\varphi_{k}$ can be modeled as a small network composed of univariate layers, and each “inner” function $\psi_{k,j}$ can similarly be enhanced with additional depth. Liu et al. [50] introduced the concept of “Kolmogorov-Arnold depth”, proposing that certain functions, which cannot be smoothly approximated with only two layers, may benefit from deeper nested structures. Their study also addresses practical challenges and offers enhancements such as adaptive grid methods [152], hybrid KAN-MLP architectures [153, 154], and multi-head activation grouping [155, 117] for improved computational performance. These developments underscore the theoretical and empirical advantages of deeper KANs. By preserving the core principle of univariate composition, multi-layer KANs offer increased modeling flexibility and computational efficiency through deeper, more expressive representations [142, 156].
Figure 2: A conceptual illustration of how MLPs and KANs work. MLPs use a fully interconnected structure that mixes all input dimensions in each layer; in contrast, KANs process each dimension with a univariate function, then combine the results via summation. This dimension-wise approach can offer a more direct path to representing high-dimensional functions. Here, $\bm{\xi}$ represents the input vector, $\bm{W}_{(l)}$ and $\bm{b}_{(l)}$ refer to the weights and biases of the $l$-th layer, respectively, $\sigma$ is a fixed activation function that can take different forms, e.g., ReLU, sigmoid, sin, tanh, [157], and $\bm{\Phi}_{l}$ is a KAN layer defined as a matrix of 1-dimensional functions.
@@
A conceptual illustration of how MLPs and KANs work is shown in Fig. 2. To understand the key differences of KANs over traditional MLPs, we compare the two architectures across four dimensions: theoretical guarantees, architectural structure, representation power, and interpretability. (i) Both KANs and MLPs are universal approximators [131, 132, 133, 50], capable of representing any continuous function on a compact domain. However, the rate of approximation and the compactness of representation may differ. KANs are grounded in the Kolmogorov–Arnold representation theorem [47, 49, 61, 62], which ensures exact representation of continuous multivariate functions via nested compositions and summations of univariate functions. Suppose each network has $\bm{\theta}$ parameters and approximates a target function $y\colon\mathbb{R}^{n}\to\mathbb{R}$, with outputs $y_{{}_{\text{KAN}}}$ and $y_{{}_{\text{MLP}}}$. Then, under suitable smoothness assumptions [158, 159], we may write,
@@
$\bigl{\|}y-y_{{}_{\text{KAN}}}\bigr{\|}_{\infty}\leq C_{{}_{\text{KAN}}}(y,n)\cdot\bm{\theta}^{-\alpha},\quad\bigl{\|}y-y_{{}_{\text{MLP}}}\bigr{\|}_{\infty}\leq C_{{}_{\text{MLP}}}(y,n)\cdot\bm{\theta}^{-\beta},$ (3)
@@
for constants $\alpha>\beta>0$. In other words, the exponent $\alpha$ that governs the approximation rate in a KAN can exceed the MLP’s exponent $\beta$, leading to more efficient parameter usage, especially in problems where each dimension has substantial structure that can be exploited by dimension-wise univariate transformations [160, 161, 162]. (ii) MLPs rely on layers of linear transformations followed by fixed pointwise nonlinearities (e.g., ReLU, sigmoid), mixing all input dimensions at each layer [133, 163]. In contrast, KANs decompose multivariate functions into combinations of learnable univariate transformations along edges (Fig. 2[50, 67]. This leads to a fundamentally different parameterization strategy that emphasizes function composition over matrix multiplication. In KANs, each edge applies a univariate transformation (e.g., radial basis functions or Chebyshev polynomials, see Section 3.4[164, 63, 65], making the architecture more aligned with the structure of the Kolmogorov–Arnold theory. (iii) The representational strength of MLPs heavily depends on their width and depth, with no explicit structural guidance for decomposing input dimensions. KANs, by design, embed a strong inductive bias based on univariate superpositions. This bias can be especially beneficial in high-dimensional settings where individual input dimensions follow structured, learnable transformations [67]. (vi) KANs provide clearer interpretability by isolating the influence of each input dimension through dedicated univariate modules. This facilitates more transparent analyses of input–output relationships, a feature less accessible in traditional MLPs, where variable mixing via weight matrices obscures individual contributions [140, 45].
@@ 3.4 Choice of Univariate Basis Functions
A central design consideration in constructing KANs is the choice of univariate basis functions used in Eq. (2). A variety of parametric forms can be used, but many studies have focused on established approximation families, such as polynomials, wavelets, or radial kernels, for their well-known theoretical and practical benefits [165, 65, 63]. Traditional KAN architectures often utilize B-spline functions [166, 167, 168, 169], though recent investigations have proposed alternative bases to improve flexibility and efficiency. For example, the Chebyshev KANs [64] leveraged Chebyshev polynomials to handle functions with high oscillations, yielding better performance on tasks such as image classification and complex function approximation. FastKAN [63] replaced B-splines with Gaussian radial basis functions for significant speedups while preserving accuracy. Similarly, Wav-KAN [65] employed wavelet expansions to capture multi-resolution patterns, leading to better interpretability for certain datasets.
Jacobi polynomials, including Legendre, Chebyshev, and others, are a well-studied class for polynomial expansions. Within this class, B-splines and Chebyshev polynomials are frequently favored in KAN implementations [50, 168, 169]. KANs originally employ basis functions in the form of splines,
@@
where $B_{i}(\xi)$ are spline basis functions defined by the grid size $g$ and polynomial degree $k$ tending to provide robust local control over the function shape, and $c_{i}$, $w_{b}$, and $w_{s}$ are coefficients learned during training, thus $\bm{\theta}=\{c_{i},\;w_{b},\;w_{s}\}$. B-splines are often robust for modeling structured data, since local changes in the knot placement or coefficients can affect only a limited portion of the domain. However, increasing $g$ to fit complex data can greatly increase the number of trainable parameters $|\bm{\theta}|_{\text{KAN}}\sim O(N_{l}N_{n}^{2}(k+g))$, where $N_{l}$ is the number of layers, and $N_{n}$ is the number of neurons per layer. A potential solution to the computational challenge involves utilizing Chebyshev polynomials, which serve as a foundational tool in approximation theory and numerical analysis, by employing them as an orthogonal basis [164, 72]. Chebyshev polynomials are defined as,
@@
while, by contrast, a trigonometric formulation can be faster but may risk numerical instability if $\xi$ approaches the bounds of [1,1][-1,1][ - 1 , 1 ] during training,
@@
which can trigger undefined values and NaN losses [170, 171, 165]. Recent works in KANs have proposed careful input normalization or additional regularization to mitigate this issue [68]. To improve stability while training, Chebyshev polynomials restrict their inputs in the range [1,1][-1,1][ - 1 , 1 ]. That is accomplished by applying the $\tanh$ after each layer [104, 172, 64], changing the corresponding forward pass, described in Fig. 2 as,
$y_{{}_{\text{cKAN}}}(\bm{\xi})=(\bm{\Phi}_{L}\circ\tanh\circ\cdots\circ\tanh\circ\leavevmode\nobreak\ \bm{\Phi}_{1}\circ\tanh)(\bm{\xi}),$ (8)
Wavelet expansions offer a powerful and flexible framework for capturing localized or piecewise-regular structures in data [65]. A univariate function can be represented as,
@@
where $d_{\ell,r}$ denotes the wavelet coefficient, representing the contribution of the basis function at scale $\ell$ and location $r$, and $\zeta_{\ell,r}(\xi)$ is a scaled and shifted version of the mother wavelet. Specifically, the wavelet basis is defined as $\zeta_{\ell,r}(\xi)=\phi\left((\xi-c_{\ell,r})/\sigma_{\ell,r}\right)$, where $c_{\ell,r}$ controls the translation and $\sigma_{\ell,r}$ governs the dilation (or spread) of the wavelet function [65, 173]. In the context of Wav-KAN [65], this parameterization enables adaptive learning of localized features in each input dimension. The complete set of trainable parameters is given by $\bm{\theta}=\left\{c_{\ell,r},\ \sigma_{\ell,r},\ d_{\ell,r}\right\},$ and the total number of parameters in Wav-KAN scales as $|\bm{\theta}|_{\text{Wav-KAN}}\sim O(3N_{l}N_{n}^{2})$, where $N_{l}$ is the number of layers and $N_{n}$ is the number of wavelet nodes per layer [65]. Common choices for the mother wavelet $\phi$ include the Morlet [174], Mexican-Hat [175], Haar [176], and Daubechies wavelets [177], each offering distinct localization properties in time and frequency domains [178, 179, 180]. Recent developments have introduced novel mother wavelet designs optimized for specific applications. For instance, Levie et al. [181] proposed a variational approach to design mother wavelets with optimally localized ambiguity functions, enhancing sparsity and reducing redundancy in continuous wavelet transforms. Additionally, Gauthier et al. [182] introduced Parametric Scattering Networks, wherein wavelet parameters such as scales and orientations are learned from data, leading to improved performance in small-sample classification tasks. The wavelet-based formulation allows KANs to model fine-grained, local variations across dimensions, which is essential for learning from data with abrupt transitions, discontinuities, or highly non-uniform patterns [183, 65].
Radial basis functions (RBFs) provide yet another avenue for univariate approximation [63]. A typical representation of a univariate function using RBFs is,
@@
where $w_{r}$ are the trainable weights, $c_{r}$ are the centers, $\sigma_{r}$ are the shape (or scale) parameters, and $\kappa(\cdot)$ is a chosen radial kernel—for instance, the Gaussian kernel $\kappa(\xi)=\exp(-\gamma\,\xi^{2})$, where $\gamma=1/\sigma^{2}$. The set of all trainable parameters is $\bm{\theta}=\left\{w_{r},\,c_{r},\,\sigma_{r}\right\},$ and the total number of parameters scales as $|\bm{\theta}|_{\text{RBF}}\sim O(3N_{l}N_{n}^{2})$. RBF expansions are known for smooth interpolation properties and ease of application to certain regression tasks [184, 185]. Several types of RBFs have been developed to address diverse approximation tasks. Gaussian RBFs are ideal for smooth function modeling [186, 187], multiquadric functions perform well on scattered data [188], inverse multiquadrics offer smooth interpolation for PDEs [189], and Wendland functions enable sparse, meshfree approximations due to their compact support [190].
@@
Two other alternative approaches for implementing univariate modules in KANs are: (i) using shallow fully connected neural networks, and (ii) combining multiple basis functions. In the first approach, each univariate transformation is represented by a small multilayer perceptron. This retains the univariate nature of the architecture while leveraging GPU-accelerated training and flexible function approximation. Recent studies have shown that this method can perform competitively on high-dimensional tasks with proper tuning [147, 191]. The second approach, exemplified by FC-KAN [192], enhances representation power by combining fundamental functions such as B-splines, wavelets, and radial basis functions. Instead of relying on a single function type, FC-KAN employs strategies like element-wise operations, polynomial expansions, and concatenation to enrich the expressiveness of univariate mappings.
@@
Each family of basis functions comes with distinct advantages and challenges. In Table 1, we summarize the key characteristics of commonly used basis functions in KANs, including their parameterization, computational complexity, and typical application scenarios. Polynomials are often fast for smooth functions but can require careful scaling or recursive evaluation. Wavelets work well for localized behaviors but demand extra decisions about wavelet types and decomposition levels. RBFs excel in smooth interpolation but require a suitable kernel choice. Shallow neural networks benefit from mature deep-learning knowledge stacks, though controlling their capacity may require significant hyperparameter searches. In practice, the choice of basis functions is guided by the problem’s structure, computational efficiency, and interpretability requirements. Ongoing research continues to refine these choices to enhance KAN performance across diverse applications [64, 63, 65].
@@
Splines $\{w_{s},w_{b},c_{i}\}$ $O(N_{l}N_{n}^{2}(k+g))$ Robust for structured data with local control; high parameter count for complex functions [167, 168, 169, 63]
Chebyshev $\{c_{i}\}$ $O(N_{l}N_{n}^{2}k)$ Efficient for smooth functions; fewer parameters and fast convergence; requires input normalization [170, 171, 68, 104, 172, 64]
Wavelets $\{d_{\ell,r},c_{\ell,r},\sigma_{\ell,r}\}$ $O(3N_{l}N_{n}^{2})$ Suitable for functions with sharp, high-amplitude oscillations; ideal for multi-resolution data [65, 173, 174, 175, 176, 177, 178, 179, 180, 181, 182]
RBFs $\{w_{r},c_{r},\sigma_{r}\}$ $O(3N_{l}N_{n}^{2})$ Effective for smooth interpolation; compact, adaptive, and easy to tune with kernel variations [184, 185, 186, 187, 188, 189, 190]

### v2

@@ 3 Kolmogorov-Arnold Networks
The exploration of the Kolmogorov-Arnold representation theorem in the context of neural networks has evolved significantly over the years. Early investigations by Sprecher & Draghici sprecher2002space and Köppen koppen2002training focused on its applicability in network training, but these efforts were largely confined to the original depth-2, width-$(2n+1)$ formulation, as presented in Eq. (1). Challenges such as the non-smoothness of the inner functions and the difficulty of efficient training limited their practical implementation. Subsequent works, including Montanelli & Yang montanelli2020error and Lai & Shen lai2021kolmogorov , have explored the theoretical implications and potential benefits of the theorem, though they did not fully exploit modern optimization techniques like backpropagation. More studies have revisited these limitations from complementary perspectives: Fakhoury et al. fakhoury2022exsplinet mitigate non-smoothness and improve empirical trainability by replacing the original inner mappings with smooth, locally supported B-spline bases (explicitly differentiable for degree greater than two); by contrast, He he2023optimal provides a constructive theoretical analysis of the optimal expressive power of ReLU deep neural networks within the Kolmogorov framework, strengthening approximation guarantees but not directly addressing smoothness or practical training-stability. Thus, while some work combines architectural choices and empirical evaluation to mitigate the classical obstacles, other contributions provide complementary theoretical foundations that justify and guide modern implementations. Building on KART and the aforementioned foundational work, Kolmogorov–Arnold Networks liu2024kan , liu2024kan2 , li2024kolmogorov , bozorgasl2405wav , ss2024chebyshev have emerged as a modern computational architecture. KANs implement the classical theorem stating that any continuous function of $n$ variables can be expressed as nested sums and compositions of univariate functions. By leveraging this decomposition, KANs construct neural network-like structures that balance expressiveness and computational efficiency, offering a promising alternative to traditional MLPs for scientific computing and high-dimensional approximation tasks ta2025af , he2024mlp . In practice, KANs can be extended beyond the classical depth-2 structure by stacking multiple layers of univariate transformations, enabling deeper architectures that support hierarchical function approximation and improved representational capacity schmidt2021kolmogorov , liu2024kan , yang2024kolmogorov , liu2024kan2 .
@@ 3.1 Two-Layer KANs
A two-layer Kolmogorov-Arnold Network girosi1989representation , ismailov2014approximation , schmidt2021kolmogorov , poluektov2023construction as expressed in Eq. (1) aims to approximate a function, $F:\mathbb{R}^{n}\to\mathbb{R},$ by representing it using a depth-2, width-$(2n+1)$ architecture analogous to the Kolmogorov-Arnold superposition, where each $\psi_{k,j}$ and $\varphi_{k}$ is a univariate, continuous (and in practice differentiable) function. In a two-layer Kolmogorov-Arnold Network: (i) The first layer consists of functions $\psi_{k,j}:\mathbb{R}\to\mathbb{R}$ for $j=1,\dots,n$ and $k=0,\dots,2n$, whose outputs are summed over $j$ (for each fixed $k$); (ii) The second layer consists of univariate functions $\varphi_{k}:\mathbb{R}\to\mathbb{R}$ that take the sum from the first layer as input, and are subsequently summed across $k=0,\dots,2n$ to produce a final scalar output. Hence, the original $n$-dimensional input is decomposed into univariate transformations, and the network’s trainable parameters correspond to the internal representations of these $\psi_{k,j}$ and $\varphi_{k}$ functions.
@@ 3.2 Multi-Layer KANs
A multi-layer KAN liu2024kan is defined as,
@@
where each $\phi_{k}^{(m)}$ remains a univariate function, but organized in a layered, nested manner. This architecture extends the classical depth-2 superposition model by allowing each univariate component to be represented as a composition of multiple univariate transformations. Consequently, each “outer” function $\varphi_{k}$ can be modeled as a small network composed of univariate layers, and each “inner” function $\psi_{k,j}$ can similarly be enhanced with additional depth. Liu et al. liu2024kan introduced the concept of “Kolmogorov-Arnold depth”, proposing that certain functions, which cannot be smoothly approximated with only two layers, may benefit from deeper nested structures. Their study also addresses practical challenges and offers enhancements such as adaptive grid methods rigas2024adaptive , hybrid KAN-MLP architectures li2024hkan , chechkin2025hybrid , and multi-head activation grouping yang2024activation , toscano2025kkans for improved computational performance. A recent theoretical advance by Kratsios et al. kratsios2025kolmogorov further reinforces the mathematical underpinnings of multi-layer KANs. They rigorously proved that residual multi-layer KANs (Res-KANs) can optimally approximate functions belonging to spaces chui1992general , devore1988interpolation , even on complex or fractal domains. Furthermore, they bounded the pseudodimension of Res-KANs to obtain dimension-independent sample complexity estimates, showing that Res-KANs can efficiently learn smooth functions even in high-dimensional spaces. These results establish a solid theoretical foundation supporting the expressive power and generalization ability of deep KAN architectures. These developments underscore the theoretical and empirical advantages of deeper KANs. By preserving the core principle of univariate composition, multi-layer KANs offer increased modeling flexibility and computational efficiency through deeper, more expressive representations liu2015kolmogorov , polar2021deep .
Figure 2: A conceptual illustration of how MLPs and KANs work. MLPs use a fully interconnected structure that mixes all input dimensions in each layer; in contrast, KANs process each dimension with a univariate function, then combine the results via summation. This dimension-wise approach can offer a more direct path to representing high-dimensional functions. Here, $\bm{\xi}$ represents the input vector, $\bm{W}_{(l)}$ and $\bm{b}_{(l)}$ refer to the weights and biases of the $l$-th layer, respectively, $\sigma$ is a fixed activation function that can take different forms, e.g., ReLU, sigmoid, sin, tanh, faroughi2023physics , and $\bm{\Phi}_{l}$ is a KAN layer defined as a matrix of 1-dimensional functions.
@@
A conceptual illustration of how MLPs and KANs work is shown in Fig. 2. To understand the key differences of KANs over traditional MLPs, we compare the two architectures across four dimensions: theoretical guarantees, architectural structure, representation power, and interpretability. (i) Both KANs and MLPs are universal approximators cybenko1989approximation , hornik1989multilayer , hornik1991approximation , liu2024kan , capable of representing any continuous function on a compact domain. However, the rate of approximation and the compactness of representation may differ. KANs are grounded in the Kolmogorov–Arnold representation theorem kolmogorov1957representations , arnold1959representation , sprecher2002space , koppen2002training , which ensures exact representation of continuous multivariate functions via nested compositions and summations of univariate functions. Let each network be parameterized by a vector of trainable parameters $\bm{\theta}$, with $|\bm{\theta}|$ denoting the total number of parameters. Moreover, consider a target function $y\colon\mathbb{R}^{n}\to\mathbb{R}$, which is approximated by the networks with outputs $y_{{}_{\text{KAN}}}$ and $y_{{}_{\text{MLP}}}$. Then, under suitable smoothness assumptions barron2002universal , devore1998nonlinear , we may write,
@@
$\bigl\|y-y_{{}_{\text{KAN}}}\bigr\|_{\infty}\leq C_{{}_{\text{KAN}}}(y,n)\cdot|\bm{\theta}|^{-\alpha},\quad\bigl\|y-y_{{}_{\text{MLP}}}\bigr\|_{\infty}\leq C_{{}_{\text{MLP}}}(y,n)\cdot|\bm{\theta}|^{-\beta},$ (3)
@@
for constants $\alpha>\beta>0$. In other words, the exponent $\alpha$ that governs the approximation rate in a KAN can exceed the MLP’s exponent $\beta$, leading to more efficient parameter usage, especially in problems where each dimension has substantial structure that can be exploited by dimension-wise univariate transformations sharma2020neural , michaud2023precision , poggio2020theoretical . (ii) MLPs rely on layers of linear transformations followed by fixed pointwise nonlinearities (e.g., ReLU, sigmoid), mixing all input dimensions at each layer hornik1991approximation , karlik2011performance . In contrast, KANs decompose multivariate functions into combinations of learnable univariate transformations along edges (Fig. 2) liu2024kan , liu2024kan2 . This leads to a fundamentally different parameterization strategy that emphasizes function composition over matrix multiplication. In KANs, each edge applies a univariate transformation (e.g., radial basis functions or Chebyshev polynomials, see Section 3.4) rivlin2020chebyshev , li2024kolmogorov , bozorgasl2405wav , making the architecture more aligned with the structure of the Kolmogorov–Arnold theory. (iii) The representational strength of MLPs heavily depends on their width and depth, with no explicit structural guidance for decomposing input dimensions. KANs, by design, embed a strong inductive bias based on univariate superpositions. This bias can be especially beneficial in high-dimensional settings where individual input dimensions follow structured, learnable transformations liu2024kan2 . (vi) KANs provide clearer interpretability by isolating the influence of each input dimension through dedicated univariate modules. This facilitates more transparent analyses of input–output relationships, a feature less accessible in traditional MLPs, where variable mixing via weight matrices obscures individual contributions kilani2024kolmogorov , cruz2025state .
@@ 3.4 Choice of Univariate Basis Functions
A central design consideration in constructing KANs is the choice of univariate basis functions used in Eq. (2). A variety of parametric forms can be used, but many studies have focused on established approximation families, such as polynomials, wavelets, or radial kernels, for their well-known theoretical and practical benefits trefethen2019approximation , bozorgasl2405wav , li2024kolmogorov . Traditional KAN architectures often utilize B-spline functions de1978practical , bohra2020learning , vaca2024kolmogorov , samadi2024smooth , though recent investigations have proposed alternative bases to improve flexibility and efficiency. For example, the Chebyshev KANs ss2024chebyshev leveraged Chebyshev polynomials to handle functions with high oscillations, yielding better performance on tasks such as image classification and complex function approximation. FastKAN li2024kolmogorov replaced B-splines with Gaussian radial basis functions for significant speedups while preserving accuracy. Similarly, Wav-KAN bozorgasl2405wav employed wavelet expansions to capture multi-resolution patterns, leading to better interpretability for certain datasets.
Jacobi polynomials, including Legendre, Chebyshev, and others, are a well-studied class for polynomial expansions. Within this class, B-splines and Chebyshev polynomials are frequently favored in KAN implementations liu2024kan , vaca2024kolmogorov , samadi2024smooth . KANs originally employ basis functions in the form of splines,
@@
where $B_{i}(\xi)$ are spline basis functions defined by the grid size $g$ and polynomial degree $k$ tending to provide robust local control over the function shape, and $c_{i}$, $w_{b}$, and $w_{s}$ are coefficients learned during training, thus $\bm{\theta}=\{c_{i},\;w_{b},\;w_{s}\}$. B-splines are often robust for modeling structured data, since local changes in the knot placement or coefficients can affect only a limited portion of the domain. However, increasing $g$ to fit complex data can greatly increase the number of trainable parameters $|\bm{\theta}|_{\text{KAN}}\sim O(N_{l}N_{n}^{2}(k+g))$, where $N_{l}$ is the number of layers, and $N_{n}$ is the number of neurons per layer. A potential solution to the computational challenge involves utilizing Chebyshev polynomials, which serve as a foundational tool in approximation theory and numerical analysis, by employing them as an orthogonal basis rivlin2020chebyshev , schmidt2021kolmogorov . Chebyshev polynomials are defined as,
@@
while, by contrast, a trigonometric formulation can be faster but may risk numerical instability if $\xi$ approaches the bounds of [1,1][-1,1] during training,
@@
which can trigger undefined values and NaN losses boyd2001chebyshev , mason2002chebyshev , trefethen2019approximation . Recent works in KANs have proposed careful input normalization or additional regularization to mitigate this issue mostajeran2025scaled . To improve stability while training, Chebyshev polynomials restrict their inputs in the range [1,1][-1,1]. That is accomplished by applying the $\tanh$ after each layer guo2024physics , hu2024tackling , ss2024chebyshev , changing the corresponding forward pass, described in Fig. 2 as,
$y_{{}_{\text{cKAN}}}(\bm{\xi})=(\bm{\Phi}_{L}\circ\tanh\circ\cdots\circ\tanh\circ\penalty 10000\ \bm{\Phi}_{1}\circ\tanh)(\bm{\xi}),$ (8)
Wavelet expansions offer a powerful and flexible framework for capturing localized or piecewise-regular structures in data bozorgasl2405wav . A univariate function can be represented as,
@@
where $d_{\ell,r}$ denotes the wavelet coefficient, representing the contribution of the basis function at scale $\ell$ and location $r$, and $\zeta_{\ell,r}(\xi)$ is a scaled and shifted version of the mother wavelet. Specifically, the wavelet basis is defined as $\zeta_{\ell,r}(\xi)=\phi\left((\xi-c_{\ell,r})/\sigma_{\ell,r}\right)$, where $c_{\ell,r}$ controls the translation and $\sigma_{\ell,r}$ governs the dilation (or spread) of the wavelet function bozorgasl2405wav , mostajeran2023novel . In the context of Wav-KAN bozorgasl2405wav , this parameterization enables adaptive learning of localized features in each input dimension. The complete set of trainable parameters is given by $\bm{\theta}=\left\{c_{\ell,r},\ \sigma_{\ell,r},\ d_{\ell,r}\right\},$ and the total number of parameters in Wav-KAN scales as $|\bm{\theta}|_{\text{Wav-KAN}}\sim O(3N_{l}N_{n}^{2})$, where $N_{l}$ is the number of layers and $N_{n}$ is the number of wavelet nodes per layer bozorgasl2405wav . Common choices for the mother wavelet $\phi$ include the Morlet bhat2024novel , Mexican-Hat ilyas2024design , Haar ramalakshmi2024identity , and Daubechies wavelets ali2024short , each offering distinct localization properties in time and frequency domains mallat1999wavelet , daubechies1992ten , torrence1998practical . Recent developments have introduced novel mother wavelet designs optimized for specific applications. For instance, Levie et al. levie2021wavelet proposed a variational approach to design mother wavelets with optimally localized ambiguity functions, enhancing sparsity and reducing redundancy in continuous wavelet transforms. Additionally, Gauthier et al. gauthier2022parametric introduced Parametric Scattering Networks, wherein wavelet parameters such as scales and orientations are learned from data, leading to improved performance in small-sample classification tasks. The wavelet-based formulation allows KANs to model fine-grained, local variations across dimensions, which is essential for learning from data with abrupt transitions, discontinuities, or highly non-uniform patterns bruna2013invariant , bozorgasl2405wav .
Radial basis functions (RBFs) provide yet another avenue for univariate approximation li2024kolmogorov . A typical representation of a univariate function using RBFs is,
@@
where $w_{r}$ are the trainable weights, $c_{r}$ are the centers, $\sigma_{r}$ are the shape (or scale) parameters, and $\kappa(\cdot)$ is a chosen radial kernel, for instance, the Gaussian kernel $\kappa(\xi)=\exp(-\gamma\,\xi^{2})$, where $\gamma=1/\sigma^{2}$. The set of all trainable parameters is $\bm{\theta}=\left\{w_{r},\,c_{r},\,\sigma_{r}\right\},$ and the total number of parameters scales as $|\bm{\theta}|_{\text{RBF}}\sim O(3N_{l}N_{n}^{2})$. RBF expansions are known for smooth interpolation properties and ease of application to certain regression tasks mostajeran2023radial , liu2025solving , muthusamy2025economic , mirzaei2024parallel . Several types of RBFs have been developed to address diverse approximation tasks. Gaussian RBFs are ideal for smooth function modeling bouzidi2025gaussian , buhmann2003radial , stenkin2024mathematical , multiquadric functions perform well on scattered data liu2025efficient , inverse multiquadrics offer smooth interpolation for PDEs ku2024deep , and Wendland functions enable sparse, meshfree approximations due to their compact support wendland2004scattered .
@@
Two other alternative approaches for implementing univariate modules in KANs are: (i) using shallow fully connected neural networks, and (ii) combining multiple basis functions. In the first approach, each univariate transformation is represented by a small multilayer perceptron. This retains the univariate nature of the architecture while leveraging GPU-accelerated training and flexible function approximation. Recent studies have shown that this method can perform competitively on high-dimensional tasks with proper tuning lai2021kolmogorov , zhang2021multiscale . The second approach, exemplified by FC-KAN ta2024fc , enhances representation power by combining fundamental functions such as B-splines, wavelets, and radial basis functions. Instead of relying on a single function type, FC-KAN employs strategies like element-wise operations, polynomial expansions, and concatenation to enrich the expressiveness of univariate mappings. A related line of work by Fakhoury & Speleers fakhoury2025expressivity analyzed the ExSpliNet model, a KAN-inspired architecture based on multivariate B-spline representations, and provided constructive approximation results demonstrating its strong expressivity and ability to mitigate the curse of dimensionality.
@@
Each family of basis functions comes with distinct advantages and challenges. In Table 1, we summarize the key characteristics of commonly used basis functions in KANs, including their parameterization, computational complexity, and typical application scenarios. Polynomials are often fast for smooth functions but can require careful scaling or recursive evaluation. Wavelets work well for localized behaviors but demand extra decisions about wavelet types and decomposition levels. RBFs excel in smooth interpolation but require a suitable kernel choice. Shallow neural networks benefit from mature deep-learning knowledge stacks, though controlling their capacity may require significant hyperparameter searches. In practice, the choice of basis functions is guided by the problem’s structure, computational efficiency, and interpretability requirements. Ongoing research continues to refine these choices to enhance KAN performance across diverse applications ss2024chebyshev , li2024kolmogorov , bozorgasl2405wav .
@@
Splines $\{w_{s},w_{b},c_{i}\}$ $O(N_{l}N_{n}^{2}(k+g))$ Robust for structured data with local control; high parameter count for complex functions bohra2020learning , vaca2024kolmogorov , samadi2024smooth , li2024kolmogorov
Chebyshev $\{c_{i}\}$ $O(N_{l}N_{n}^{2}k)$ Efficient for smooth functions; fewer parameters and fast convergence; requires input normalization boyd2001chebyshev , mason2002chebyshev , mostajeran2025scaled , guo2024physics , hu2024tackling , ss2024chebyshev
Wavelets $\{d_{\ell,r},c_{\ell,r},\sigma_{\ell,r}\}$ $O(3N_{l}N_{n}^{2})$ Suitable for functions with sharp, high-amplitude oscillations; ideal for multi-resolution data bozorgasl2405wav , mostajeran2023novel , bhat2024novel , ilyas2024design , ramalakshmi2024identity , ali2024short , mallat1999wavelet , daubechies1992ten , torrence1998practical , levie2021wavelet , gauthier2022parametric
RBFs $\{w_{r},c_{r},\sigma_{r}\}$ $O(3N_{l}N_{n}^{2})$ Effective for smooth interpolation; compact, adaptive, and easy to tune with kernel variations mostajeran2023radial , liu2025solving , bouzidi2025gaussian , buhmann2003radial , liu2025efficient , ku2024deep , wendland2004scattered , muthusamy2025economic , mirzaei2024parallel , stenkin2024mathematical